OpenStudy (anonymous):

What are the roots of 4x2 + 176 = 0?

OpenStudy (solomonzelman):

\[4x^2+176=0,~~~~~~factor~~out~~of~~4,\]\[4(x^2+44)=0\]\[x^2+44=0\]\[x^2=-44\]\[x=~±~2i \times \sqrt{11}\]

OpenStudy (solomonzelman):

are you familiar with imaginary numbers, if yes, there it is, if no, state no solution.